Let X be a random variable such that -1, with probability q 1 with probability 1-q,

a) Find the moment generating function (mgf) of X.

b) Using part a), that is the mgf of X, find the expected value (E[X]) and (V ar[X]).
Answer #1

a)

as we know that mgf=Mx(t)=\sum etxP(x)

Mx(t) =e-t*q+et*(1-q)

b)

first derivation of mgf :M'x(t)=(d/dt)*(e-t*p+et*(1-p))=-qe-t+(1-q)et

hence E(X)=M'x(0)=-qe-0+(1-q)e0 =-q+1-q =1-2q

second derivation of mgf :M''x(t)=(d/dt)*(-qe-t+(1-q)et )=qe-t+(1-q)et

E(X2)=:M''x(0)=qe-0+(1-q)e0 =q+1-q =1

therefore Var(X)=E(X2)-(E(X))2 =1-(1-2q)2 =4q-4q2 =4q(1-q)